### v1

# Mamo: a Mathematical Modeling Benchmark with Solvers

Abstract
Mathematical modeling involves representing real-world phenomena, systems, or problems using mathematical expressions and equations to analyze, understand, and predict their behavior. Given that this process typically requires experienced experts, there is an interest in exploring whether Large Language Models (LLMs) can undertake mathematical modeling to potentially decrease human labor. To evaluate of LLMs in mathematical modeling, we introduce a new benchmark, Mamo, that transcends traditional result-oriented assessments. Unlike conventional methods that primarily assess LLMs based on the accuracy of solutions to mathematical problems, our approach offers deeper insight into the modeling process itself. By focusing on the processes LLMs undertake rather than the correctness of their final solutions, Mamo pioneers a novel evaluation paradigm. This shift underscores the importance of understanding the inherent modeling capabilities of LLMs, paving the way for a more nuanced and comprehensive analysis of their problem-solving strategies. Our work marks a significant advancement in the field, suggesting a new direction for future research by emphasizing the evaluation of LLMs’ modeling processes over the mere correctness of answers. This benchmark not only facilitates a better understanding of LLMs’ mathematical modeling capabilities but also sets a new standard for evaluating their performance in complex problem-solving scenarios.
1 Introduction
Recent advancements in Large Language Models (LLMs) have garnered widespread interest, demonstrating remarkable capabilities across a broad spectrum of natural language processing tasks Ouyang et al. (2022); OpenAI (2023); Nijkamp et al. (2022); Tang et al. (2024). However, the domain of mathematics remains a critical aspect of LLM evaluation Wei et al. (2022). This focus not only gauges LLMs’ proficiency in comprehending and addressing mathematical challenges but also serves as a profound indicator of their underlying abilities in abstract conceptualization and logical reasoning, among other intellectual faculties.
Natural language captures human complexity and nuance, while mathematical language excels in precision and modeling the natural world (Giordano et al., 2013). Bridging their gap requires high-level cognitive skills akin to Artificial General Intelligence (AGI), a task at which LLMs show promise. However, the challenge lies in the evaluation of modeling, as perfectly representing reality may be inherently elusive.
In this endeavor, solvers [^1] In optimization, solvers are algorithms used to find the best solution to a problem. They search through possible options to maximize or minimize an objective under given constraints if any. By using solves, The challenge becomes problem formulation under a given solver or a solver set. becomes critically important as it standardizes the modeling output and could provide a way to falsifiably validate the produced answers. Recent studies have already combined LLMs with solvers Yang et al. (2023); Feng et al. (2023); Pan et al. (2023). In the field of optimization, AhmadiTeshnizi et al. (2023b) introduced OptiMUS, an LLM-based agent that uses a solver to address optimization problems. The key task for LLM in OptiMUS involves first generating a mathematical formulation of a real-world problem, followed by writing Python code to invoke a solver. This implies the potential for testing LLM’s modeling capabilities.
We thus develope Mamo, a new benchmark that leverages solvers to assess the mathematical modeling prowess of large language models (LLMs). Our approach shifts from conventional outcome-focused evaluation to process-focused evaluation, offering in-depth insights into the LLMs’ problem-solving strategies. We concentrate on the models’ mathematical modeling skills, delegating the task of solving abstract problems to solvers to circumvent errors arising from the models’ computational capabilities.
The contributions of this paper are as below:
  1. 1.
    Clarification of a Mathematical Modeling Framework for LLMs in Natural Language Domains: We have meticulously clarified the concept of mathematical modeling within Large Language Models (LLM), focusing on the realm of natural language challenges. This effort has culminated in the establishment of a comprehensive framework designed to guide the mathematical modeling process in LLMs, addressing both theoretical underpinnings and practical applications.
  2. 2.
    Innovation in Benchmarking Strategy through Solver Integration: Our research introduces a pioneering benchmarking strategy for the assessment of mathematical modeling process. By incorporating solvers into the benchmarking methodology, this strategy facilitates a rigorous evaluation process. This novel approach sets a new standard for the assessment in NLP research.
  3. 3.
    Creation of the Mamo Benchmark for Extensive Mathematical Modeling Assessment: We have developed a new benchmark, named Mamo, specifically tailored for the evaluation of mathematical modeling capabilities. Mamo encompasses a wide array of modeling questions, focusing on areas such as ordinary differential equations and optimization problems within linear programming and mixed-integer linear programming frameworks. With a total of 1059 meticulously curated questions, Mamo stands as a comprehensive tool for assessing the breadth and depth of mathematical modeling proficiency.
2 Related Work
Recent studies have made significant strides in applying LLMs to mathematical problem-solving, introducing innovative datasets for comprehensive evaluation. Frieder et al. (2024) and Yuan et al. (2023) have focused on datasets that test LLMs across various math problems and arithmetic expressions, respectively. Lewkowycz et al. (2022) explored training models using natural language paired with LaTeX-formatted mathematical content from arXiv, enhancing syntactic understanding. Zhang et al. (2024) developed the CARP dataset for computation-intensive challenges, while He et al. (2024) introduced OlympiadBench, a bilingual and multimodal benchmark for competition-level mathematical reasoning and proofs. Liu et al. (2024) constructed MathBench, which is designed with a carefully structured difficulty hierarchy and focuses on assessing the understanding of theoretical knowledge in LLMs. AhmadiTeshnizi et al. (2023a) constructed NLP4LP, a benchmark comprising 52 linear programming (LP) and mixed-integer linear programming (MILP) problems. These efforts collectively push the boundaries of LLMs in mathematical cognition, showcasing the growing complexity and specificity of tasks LLMs are being trained to tackle.
The pursuit of enhancing and devising novel solvers has seen a variety of research efforts, with a notable focus on the integration of LLMs. Yang et al. (2023) investigated the potential of employing LLMs as direct solvers, pioneering the exploration of LLMs’ autonomous problem-solving capabilities. He-Yueya et al. (2023) introduced a hybrid approach that marries an LLM with an external symbolic solver, specifically for equation solving, marking a significant advancement in combining neural and symbolic computation. Furthermore, Pan et al. (2023) developed LOGIC-LM, a cutting-edge framework that fuses LLMs with symbolic solvers, aimed at enhancing the solving of logical problems. In a closely related study, AhmadiTeshnizi et al. (2023a) introduced OptiMUS, an LLM-based agent explicitly crafted for formulating and solving optimization problems, thus demonstrating the broad applicability and potential of LLMs in tackling intricate computational challenges. These endeavors underscore the innovative directions being explored in the realm of solver development, leveraging the robust capabilities of LLMs.
3 Background
3.1 Mathematical Modeling
Giordano et al. (2013) emphasize the role of mathematical models as intermediaries that transform real-world problems into structured mathematical forms, facilitating the discovery of solutions. Despite LLMs’ proven competencies across various tasks, their capacity for developing mathematical models remains uncertain. This involves two key phases: model formulation, which demands a thorough comprehension of the problem, and model resolution, typically addressed by computational solvers. To accurately gauge LLMs’ potential in this domain, our strategy focuses on their ability to construct mathematical models, entrusting calculations to specialized solvers. This approach highlights LLMs’ creativity and generalization skills, simultaneously capitalizing on solvers’ computational efficiency.
Motivations to benchmark mathematical modeling for LLMs
We evaluate mathematical models for two reasons. The first reason is out of curiosity: We aim to investigate whether LLMs can do what expert humans can. Mathematical modeling is considered a highly advanced capability that we believe experts possess. This is a beneficial test for exploring the boundaries of LLMs’ capabilities and to what extent they can approach human intelligence. On the other hand, in terms of evaluating the models themselves, most current benchmarks, especially those related to mathematics like GSM8K Cobbe et al. (2021), etc., show that the performance has become saturated. Even a 7B model Yu et al. (2023b); Liu et al. (2023) (sometimes with a verifier-like model Yu et al. (2023a)) can achieve a accuracy exceeding 80%, closely approaching that of GPT 4. Therefore, we believe there is a need to construct more challenging mathematical datasets. Traditional mathematical modeling is a labor-intensive, custom process reliant on human expertise, raising the question: Can large language models automate or assist in this process?
3.2 Using Solvers in Mathematical Modeling
The model generated by the LLM is then subjected to verification against real-world data to assess its validity, ensuring it addresses the problem logically and sensibly. The solver, an algorithmic tool designed to find solutions to mathematical models, is also involved in this process and it is also beneficial for verification.
Different types of solvers are specialized for various classes of problems, such as optimization solvers. However, they share the common property: They are all algorithmic tools designed to find solutions to mathematical models.
Rationale of Using solver in Mathematical Modeling for LLMs These solvers operationalize the abstract constructs created by LLMs, translating them into tangible outputs that can be analyzed and validated against real-world data and phenomena. The solver is a crucial instrument that resolves the model presented by the LLM, and by comparing the solution it provides with the correct answer, we assess the accuracy of the mathematical model generated by the LLM.
4 Philosophy of Mamo Benchmark
4.1 Methodology: exact answer verification with solvers
We utilize solver to resolve this mathematical model. The solver operates at the final state of the question-solving process, meaning it directly computes the final answer from the model without intermediate steps. By executing the solver and juxtaposing its output ($\widehat{A}$) with the correct answer (A) to the original problem, we can verify the accuracy of the Language Model’s mathematical model(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$). This final-state approach ensures that our assessment is squarely focused on the Language Model’s capacity to model the problem, abstracting from any additional computational steps that might otherwise obscure the evaluation of the model’s validity.
Figure 1: The pipeline to use exact answer verification via an additional solver.
4.2 Scope of Mathematical Modeling
We construct the category of Mathematical Modeling by referring to Giordano et al. (2013). In the context of Natural Language we propose a framework in mathematical modelling in LLM, specified in solving natural language problem, see Tab. 1.
Category In Mamo
Changes and Differences
Proportionality and Geometric Similarity
Optimization Problem Modeling
Differential Equations
Probabilistic Modeling
Table 1: The taxonomy of mathematical modeling. In Mamo benchmark, we only include the categories where sophisticated Solvers is available.
Given the broad spectrum of mathematical modeling and the availability of specific solvers, our benchmark is meticulously designed to encompass domains where these tools can be effectively utilized. This section outlines the rationale for our focus on certain areas of mathematical modeling, based on solver availability and the clarity of the modeling process.
Availability of Sophisticated Solvers: Advanced optimization solvers, such as COPT Ge et al. (2023), along with Python libraries for solving ordinary differential equations (ODEs), offer a strong foundation for testing the capabilities of Large Language Models (LLMs). These tools enable a detailed and automated evaluation of the LLMs’ abilities to abstract, formulate, and accurately solve mathematical problems within these specific domains.
Challenges in Modeling Changes and Differences, Proportionality and Geometric Similarity: Although simpler tools like calculators could address problems related to modeling changes and differences, proportionality and geometric similarity, the lack of a defined, structured modeling process complicates the implementation of our final-state-approach benchmark. Our objective is to assess the LLMs’ skill in developing comprehensive mathematical models that adhere to a solvable framework. The absence of a distinct modeling procedure in these areas impedes a standardized and objective assessment of an LLM’s modeling efficiency.
Following the considerations mentioned above, we have developed our benchmark in the fields of optimization, and ODE, due to the lack of advanced and universal Partial Differential Equations (PDE) solvers. In the optimization segment, our focus is on Linear Programming (LP) and Mixed-Integer Linear Programming (MILP) problems. Compared to NLP4LP, introduced by AhmadiTeshnizi et al. (2023b) and focusing exclusively on optimization, our benchmark includes a larger number of problems. This benchmark set is highly scalable, with potential future expansions into nonlinear optimization and probabilistic modeling. Additionally, as new and advanced solvers emerge, it could be expanded into areas like Partial Differential Equations (PDE) and beyond.
5 Data Collection and Synthesis
5.1 Data Selection and Source Credibility
The data set for our benchmark consists of a blend of manually selected and GPT-generated questions. All items within the data set have been subjected to a meticulous review process conducted by qualified collectors and domain experts to ensure their validity and relevance.
Annotator Qualifications
The benchmarking process is conducted by a team of qualified individuals with extensive expertise in various mathematical disciplines.  [^2] See the details of the annotator qualifications in Appendix. B:
5.1.1 Manually Selected Questions
  • Source and Reliability: When doing manually selecting, we refer to the textbooks, exercises and the solutions in ODE and optimization area(Thomas et al., 2016; Stewart et al., 2014; Lial et al., 2017; Braun, 1993; Boyce & DiPrima, 2012; Giordano et al., 2014; Zill, 2013; Giordano et al., 2014; Bertsimas & Tsitsiklis, 1997; Hurlbert, 2010). The selecting process focus on deriving the mathematical model in the questions and assign a reasonable scenario. The source ensure the correctness of the answer of each question.
  • Rewriting Process: Each question is meticulously rewritten to fit the benchmark’s criteria(the properties in Section 5.3), which may involve assigning new scenarios or adapting the question and its answer to align with our testing methodology. This process also circumvents potential licensing issues, as the original content is significantly transformed.
Data Synthesis
The process of annotating data starts with the creation of typical mathematical constructs, into which random parameters are introduced to give them specificity. This is followed by the computation of answers to establish a ground truth. Subsequently, GPT-4 is employed to craft real-life scenarios based on these predefined models, effectively translating abstract mathematical concepts into tangible, context-rich problems. This approach not only tests the model’s ability to handle mathematical reasoning but also its capacity to contextualize mathematical principles within real-world situations. See the example in Appendix. E.
Cross Validation
GPT-4 is employed to confirm that the newly generated questions are amenable to being solved by solvers using the final-state approach, ensuring the benchmark’s focus remains on the LLM’s modeling capabilities. Moreover, we employ a comprehensive review process, re-examining each question for consistency and correctness, thereby ensuring the quality of our dataset. See the review details in Appendix. A.
5.2 Data Statistics and Visualization
In the section on Ordinary Differential Equations (ODE), we present a total of 346 problems: 196 are based on first-order equations, 110 on second-order equations, and 40 on systems of equations, offering a comprehensive exploration across different levels.
The optimization section is divided into two segments: Easy LP, featuring 652 high school-level Mixed Integer Linear Programming (MILP) problems, and Complex LP, comprising 211 undergraduate-level problems that integrate both LP and MILP.
See the word cloud of the combined data in Figure 2
Figure 2: The word cloud of the data after filtering the requirement words such as ’significant figures’, ’rounded’
5.3 Guidance of Data Synthesis
The synthesis of problems for our benchmark are governed by specific criteria designed to evaluate the mathematical modeling capacity of Large Language Models (LLMs). The guidelines ensure that problems are appropriate for our final-state approach and yield numerical answers suitable for automated verification. The following are the key considerations in our benchmark guidance:
Property 1.
Final-State Approach Solvability: Problems should be structured so that they can be solved by a solver in a final-state approach, meaning the solution process should not require intermediate steps once the mathematical model has been formulated by the LLM.
Property 2.
Numerical Answers: The answers to all questions must be numerical, allowing for clear-cut evaluation and comparison with the solver’s output.
Property 3.
Significant Figures and Precision: Each question will explicitly state the required significant figures or the level of numerical precision for the answer. This clarity ensures that the LLM’s output can be properly evaluated against the solver’s results, maintaining consistency and rigor in the assessment of the model’s accuracy.
Property 4.
Modeling Capacity Test: To specifically test the LLM’s modeling capabilities:
  • For Ordinary Differential Equations (ODEs), questions should focus on determining the value of the original function at a specific time $t_{0}$.
  • For optimization problems, questions should aim to find the optimal value.
These constraints facilitate the testing accuracy and the convenience of verifying potentially multiple valid solutions to a single problem.
Property 5.
Real-World Problem Context: Questions should be framed as real-world problems without explicitly stating the underlying mathematical model, challenging the LLM to demonstrate its capacity for abstraction and application in practical scenarios.
6 Evaluation
6.1 Evaluation Protocol
6.1.1 Evaluation Details
Our evaluation protocol is meticulously designed to test the LLMs’ mathematical modeling process. For ODE problems, the LLM is prompted to generate Python code from a natural language description. The code is executed, and its output is compared with the correct answer to assess accuracy. In optimization problems, the LLM is asked to express the model in .lp format, after which the COPT solver is used to find the optimal value for comparison with the correct answer.
Model output
In ODE parts, we ask LLMs to output Python code calling solvers ( solve_ivp odeint in SciPy, and dsolve in NumPy). In the optimization sections, we prompt LLMs to generate the standard .lp file. Compared to the testing methodology in NLP4LP (AhmadiTeshnizi et al., 2023b), which involves prompting the model to output Python code for invoking the solver, the .lp format more closely aligns with the standard optimization form, making it more readable. Furthermore, optimization solvers such as COPT (Ge et al., 2023) can directly read and solve the .lp file, minimizing potential issues related to the LLM’s coding capabilities. To better adhere to the format, we conducted tests using 3-shot learning. Detailed information about the number of few-shot samples used in our experiments and the result can be found in the Appendix. (see Appendix C and Appendix. G).
Mitigation of Formatting Errors
It is essential to distinguish errors in modeling from those in coding or file formatting. Therefore, if the LLM’s output contains syntax errors or incorrect .lp formatting, we need to rectify these issues without altering the underlying logic of the code or model. This step ensures that our evaluation focuses solely on the modeling process. To achieve this, we use a language model to fix any syntax errors, referring to it as the code modifier. In our experiments, we use the original language model and GPT-4-0613 as code modifiers. For the sake of reproducibility, we choose the orignal LLM; To ensure the comparison focuses exclusively on modeling capacity, we choose GPT-4-0613 as the code modifier due to its exceptional coding proficiency and operational reliability, as well as its stability than other provisional release Language Models. See the comparison of the result from different code modifier in Section 6.3 and the prompts in Appendix.  G. Besides, at the beginning of the testing, we do the first clean of the codes, clearing up the string like " “‘python “‘ " that will affect our testing.
6.1.2 Evaluation Metrics
Given the potential for minor discrepancies in numerical solutions, our protocol includes a comparison rule that accommodates for slight inaccuracies. Let $O$ denote the output from the LLM, and $A$ denote the standard answer. The comparison is defined by the following criteria  333 See the evaluation script in Appendix. F
Precision Adjustment for Floating-Point Numbers If the standard answer, $A$, is a floating-point number, the precision adjustment involves calculating the number of significant digits, $n$, after the decimal point. Both $O$ and $A$ are then scaled by $10^{n}$ for comparison.
General Comparison Criterion The overall correctness is determined based on the following conditions, applicable to both adjusted and original values:
$\text{Correctness}=\begin{cases}1&\text{if }\left|\frac{O-A}{A}\right|\leq 1\times 10^{-4}\text{ or }|O-A|\leq 1,\\ 0&\text{otherwise.}\end{cases}$ (1)
This structured evaluation ensures that the LLM’s output is accurately assessed against the standard answer with consideration for the specific requirements of numerical precision and the inherent variability in computational answers.
6.2 Benchmarking Results
Models ODE Optimization
First order Second order System All Easy LP Complex LP
GPT-4o 69.39% 73.64% 42.50% 67.63% 87.27% 23.22%
GPT-4o 69.90% 74.55% 42.50% 68.21% 87.42% 32.23%
GPT-4-turbo 66.84% 69.09% 42.50% 65.12% 87.88% 23.70%
GPT-4-turbo 67.86% 70.00% 42.50% 65.99% 87.88% 30.33%
GPT-4 60.71% 65.45% 42.50% 60.12% 86.50% 22.27%
GPT-4 61.73% 67.27% 45.00% 61.56% 86.96% 24.64%
GPT-3.5-turbo 17.35% 16.36% 7.50% 15.90% 81.29% 9.95%
GPT-3.5-turbo 25.51% 31.82% 15.00% 26.30% 84.82% 9.95%
Claude-3-Sonnet 53.57% 61.82% 30.00% 53.47% 83.59% 25.12%
Claude-3-Sonnet 55.61% 63.64% 35.00% 55.78% 84.20% 25.59%
Claude-3-Haiku 42.35% 60.91% 20.00% 45.66% 78.53% 19.91%
Claude-3-Haiku 52.04% 63.64% 22.50% 52.31% 85.89% 21.33%
Gemini-1-pro 26.53% 35.45% 2.50% 26.59% 78.99% 15.31%
Gemini-1-pro 35.71% 39.09% 7.50% 33.53% 80.06% 15.31%
Gemini-1.5-pro 66.33% 66.36% 47.50% 64.16% 85.28% 30.81%
Gemini-1.5-pro 67.35% 66.36% 52.50% 65.32% 85.28% -
DeepSeek-v2 40.82% 45.45% 32.50% 41.33 % 86.50% 6.16%
DeepSeek-v2 42.35% 64.55% 32.50% 48.27% 86.66% 6.16%
DeepSeek-coder 32.14% 49.09% 15.00% 35.55% 75.77% 18.96%
DeepSeek-coder 34.18% 49.09% 15.00% 36.71% 76.99% 18.96%
Table 2: Evaluation result on Mamo Benchmark. The remark $\dagger$ indicates it additionally uses the original LLM to rectify syntax error of the code and .lp file generated by the corresponding LLM. Here GPT-4-turbo denotes GPT-4-turbo-2024-04-09 and GPT-4o denotes GPT-4o-2024-05-13, while GPT-4 refers to GPT-4-0613
Our evaluation includes both closed-source LLMs and open-source LLMs. The closed-source commercial LLMs include the GPT-4 series OpenAI (2023), the Claude series  [^4] https://www.anthropic.com/news/claude-3-family, and the Gemini series  [^5] https://deepmind.google/technologies/gemini/. The open-source LLMs include DeepSeek-v2 DeepSeek-AI (2024) and DeepSeek-coder Guo et al. (2024). Table 2 presents the evaluation results of different language models on the Mamo Benchmark. The performance of these models is assessed both in their original form and when enhanced with the language model itself to rectify syntax errors (indicated by the dagger symbol ). Table 3 displays the performance results when GPT-4-0613 is selected as the code modifier.
Models ODE Optimization
First order Second order System All Easy LP Complex LP
GPT-4o †† 69.90% 74.55% 42.50% 68.21% 87.42% 29.86%
GPT-4-turbo †† 67.85% 70.00% 42.50% 65.99% 88.19% 28.44%
GPT-4 †† 61.73% 67.27% 45.00% 61.56% 86.96% 24.64%
GPT-3.5-turbo †† 47.96% 60.91% 20.00% 48.84% 85.43% 10.43%
Claude-3-Sonnet †† 58.16% 66.36% 35.00% 59.09% 85.12% 28.44%
Claude-3-Haiku †† 53.06% 65.45% 25.00% 53.76% 86.81% 21.33%
Gemini-1-pro †† 43.37% 54.55% 15.00% 43.64% 80.37% 16.27%
Gemini-1.5-pro †† 67.35 % 67.27% 52.50% 65.61% 85.28% 34.12%
DeepSeek-v2 †† 56.12% 68.18% 32.50% 57.23% 86.81% 6.16%
DeepSeek-coder †† 33.67% 49.09% 17.50% 36.71% 78.83% 18.96%
[…]
The evaluation results on the Mamo Benchmark reveal that GPT-4o stands out among the models in both ODE and LP tasks. Specifically, GPT-4o achieves the highest performance in ODE tasks with 69.90% in first-order and 74.55% in second-order equations, and also excels in LP tasks with 87.42% in easy LP and 32.23% in complex LP scenarios when using the LLM to rectify syntax errors. Additionally, DeepSeek-v2, as an open-source LLM, demonstrates significant capability with 48.27% overall in ODE tasks and 86.66% in easy LP tasks when enhanced with syntax correction, highlighting its potential despite being open-source.
The inclusion of the LLM itself as code modifier to rectify syntax errors results in significant improvements across various models. For instance, GPT-3.5 Turbo’s performance in ODE tasks increased from 15.90% to 26.30% overall and Claude-3-Haiku shows a remarkable improvement in ODE tasks from 45.66% to 52.31%. Similarly, DeepSeek-v2, as an open-source LLM, demonstrates significant capability with its overall ODE performance increasing from 41.33% to 48.27% when enhanced with syntax correction.
Selecting GPT-4 as a code modifier generally results in better performance than using the LLM itself for self-improvement. For instance, DeepSeek-v2’s accuracy in ODE tasks improves from 48.27% with self-improvement to 57.23% when GPT-4 is used as the code modifier. This pattern indicates that GPT-4 is a more effective code modifier, enhancing the accuracy and performance of various models more reliably than self-improvement.
6.3 On the format errors and code modifiers
Models ODE Optimization
format raw self-improved GPT-4-improved raw self-improved GPT-4-improved
GPT-4o 96.82% 99.42% 99.42% 92.93% 97.91% 97.91%
GPT-4-turbo 96.82% 99.13% 98.84% 93.86% 97.80% 97.56%
GPT-4 94.51% 97.69% 97.69% 90.96% 95.13% 95.13%
GPT-3.5-turbo 32.37% 60.40% 94.22% 87.95% 89.80% 96.76%
Claude-3-Sonnet 84.68% 93.06% 97.40% 91.89% 93.97% 96.76%
Claude-3-Haiku 78.03% 92.77% 98.27% 82.61% 92.58% 96.29%
Gemini-1.0-pro 56.36% 69.94% 88.73% 88.06% 89.11% 94.21%
Gemini-1.5-pro 93.64% 98.27% 98.84% 95.48% - 96.87%
deepseek-v2 70.81% 83.81% 96.24% 90.27% 92.00% 94.90%
deepseek-coder 88.44% 93.64% 95.95% 92.58% 96.29% 98.84%
Table 4: Correct format rate on raw, self-improved, and GPT4-improved format.
The data presented in Table 4 provides a comparative analysis of the performance (rate of correct format) of various models using different improvement methods. The table considers three scenarios: no improvement, self-improvement, and improvement provided by GPT-4.
The GPT-3.5-turbo model shows the most dramatic improvement when enhanced by GPT-4, jumping from a 32.37% success rate to 94.22%. Interestingly, Claude-3 models (Sonnet, and Haiku) and DeepSeek models (DeepSeek-v2, DeepSeek-coder) also exhibit high improvement rates when enhanced by GPT-4, highlighting the compatibility of GPT-4’s improvement methods with other models. In the realm of Optimization tasks, the GPT-3.5-turbo model again shows a significant improvement when enhanced by GPT-4, jumping from a 87.95% success rate to 96.76%.
The improvement from GPT-4 is generally superior to both self-improvement and no improvement. This indicates that GPT-4 can serve as a more effective and fair code modifier. By using GPT-4 for syntax error correction, we can better examine the modeling ability of different LLMs without the confounding factor of formatting errors. Notably, in ODE tasks, the correct format rate for all models, except one, is generally greater than 94%, with the lowest being 88.73%. In Optimization tasks, the correct format rate is even higher, with the lowest rate being 94.21%.
7 Conclusion and Future Directions
In this article, we introduce Mamo, a novel benchmark that utilizes solvers to evaluate the mathematical modeling capabilities of large language models (LLMs). We propose a shift in the evaluation paradigm from traditional outcome-oriented evaluation to process-oriented evaluation. This shift allows us to scrutinize the modeling process undertaken by LLMs, providing a more detailed and comprehensive assessment of their mathematical problem-solving abilities. Our evaluation focuses more on the mathematical modeling capabilities of the models, leaving the process of solving abstract mathematical problems to the solvers, thereby avoiding the impact of errors such as computational capabilities of the models on their actual modeling abilities.
Our benchmark represents an important step in understanding the fundamental modeling capabilities of LLMs. By focusing on the modeling process rather than the correctness of the final solution, we are able to gain valuable insights into the strengths and limitations of current LLMs. This approach provides new directions for future research, emphasizing the importance of the modeling process in solving mathematical problems.
The introduction of Mamo paves new avenues for future research. The benchmark can be expanded to include a wider variety of mathematical problems and solvers, broadening its applicability and providing a more comprehensive evaluation of the mathematical modeling capabilities of LLMs. Insights gained from the benchmark can be used to guide the development of future LLMs. By understanding the strengths and weaknesses of current models in solving mathematical problems, researchers can design new models that better handle these tasks. We will extend our model list to other models (Llama2 series, Mistral series etc) in our further study.
Limitations
While the current benchmarking methodology necessitates LLMs to generate Python code for ODEs and .lp files for optimization problems, it opens up exciting avenues for future research. The process, while effective, may inadvertently influence the LLMs’ focus towards formalization over the conceptual modeling, potentially affecting the depth of mathematical reasoning. This presents an opportunity to explore innovative benchmark designs that can better distinguish between an LLM’s modeling prowess and its formalization abilities. Future research could aim to develop methodologies that allow LLMs to demonstrate their mathematical modeling capabilities without the intermediary step of formalization, leading to a more nuanced and comprehensive assessment of their abilities
Ethics Statement:
The benchmark is designed with a stringent ethical framework to ensure that all problems are socially responsible and do not perpetuate or encourage harmful biases or stereotypes. Questions are meticulously reviewed to avoid any content that could lead to the dissemination of misinformation, support unethical practices, or cause harm to individuals or groups. This includes but is not limited to issues of privacy, security, and the potential for misuse of the model. Furthermore, the benchmark abstains from any problem that could indirectly endorse unethical behaviors or decisions in real-world scenarios, upholding the highest standards of academic integrity and social responsibility.
Appendix A Cross Review
In ODE part, we have the following review process:
  1. 1.
    We employed GPT-4 to confirm that the newly generated questions are amenable to being solved by solvers using the final-state approach. There were only 10 out of 383 questions are invalid. All were filtered, 373 questions were left.
  2. 2.
    Then we conducted a thorough evaluation of the questions. Our dataset initially comprised $373$ questions. We identified $29$ questions that lacked sufficient information, $21$ that had incorrect answers, and $18$ that featured unclear statements. Following our assessment, $27$ questions were deleted, and $41$ were corrected to ensure clarity and correctness.
Finally, we have 346 problems in ODE part.
We have also conducted the cross-review process. See the details in Appendix.  A.1 and the results on Appendix.  A.1.1.
In the optimization section, particularly with easy LP part, out of 688 entries, 12 were found to be completely invalid (not meeting the criteria), 26 had incorrect answers, and 6 had misleading descriptions. All invalid entries were filtered out, and the remaining issues were corrected manually. Among them, 8 of them were corrected and others (36 questions) were filtered out. In complex LP part, for the original 211 questions, 40 were found synthetized by wrong mathematical models. All of them were corrected.
A.1 Cross-Review Process
Furthermore, adhering to the scaling law evident in the dataset’s distribution, we selected 50 questions from the ODE parts in our dataset for a deeper analysis. Four independent reviewers (see their qualification in Section B) were tasked with assessing 50 questions each. The responses could either be numeric or labeled as “error” if a question was deemed problematic. This section outlines the metrics used to evaluate the reviewers’ responses against pre-defined correct answers, which also include the “error” label for any flawed questions. We use the same metric as described in Subsection 6.1.2. Here are our results:
A.1.1 Results
The effectiveness of the review process was quantified using the metrics defined in Subsection 6.1.2, focusing on the accuracy and inter-rater reliability among reviewers. The statistical outcomes are summarized as follows:
  • Average Cohen’s Kappa: The average Cohen’s Kappa across all reviewer pairs was 0.60, indicating a moderate to substantial agreement. This suggests that reviewers generally agreed on the classification of answers, though there were variations in some cases.
  • Minimum and Maximum Cohen’s Kappa: The minimum Kappa value recorded was 0.50, and the maximum was 0.71. The spread in these values highlights areas where alignment and training could potentially enhance consistency.
  • Accuracy Rates: The individual accuracy rates were as follows:
    • Reviewer 1: 90.0%
    • Reviewer 2: 84.0%
    • Reviewer 3: 88.0%
    • Reviewer 4: 74.0%
    These rates reflect the precision with which each reviewer matched the standard answers, including their recognition of problematic questions correctly labeled as “error.”
These results provide insights into the overall effectiveness and areas of improvement for the cross-review process, particularly in terms of aligning understanding and interpretation of the evaluation criteria among reviewers.
Appendix B Annotator Qualifications
The benchmarking process benefits from the expertise of reviewers with robust backgrounds in mathematics, underscored by their academic accomplishments. The team is composed of individuals whose education encompasses critical mathematical disciplines, including calculus (I and II), linear algebra (both introductory and advanced levels), optimization, probability, and ordinary differential equations. The average Grade Point Average (GPA) across these foundational courses is approximately 3.89 out of a 4.0 scale. This high level of academic achievement demonstrates the reviewers’ strong grasp of essential mathematical concepts and their application. Furthermore, the team is enriched by members who hold Ph.D. degrees in mathematics, further solidifying the depth of expertise and analytical skills available for the benchmarking process.
B.1 Qualification of Reviewers
We select the reviewers mentioned in Appendix  A.1 based on the following criteria:
  1. 1.
    Undergraduate students who have taken fundamental mathematical courses (Calculus and Linear Algebra) and have an average GPA of at least 3.5/4.0.
  2. 2.
    Have completed or are currently taking courses in Ordinary Differential Equations (ODE).
  3. 3.
    Meet one of the following English proficiency requirements: TOEFL > 80, IELTS > 6.5, a minimum grade of B+ in all English classes, or a Gaokao English score > 120.
These criteria ensure their ability to read questions in English and solve ODE problems.
Appendix C Few-shot Experiment
To test the sensitivity of different models to the number of prompts, we conducted a sensitivity analysis using the scaling law to examine the effect of the number of shots on model performance. In our analysis, we tested the performances of 11 models on ODE problems using prompts with 0 shot, 1 shot, 3 shots, 5 shots, and 10 shots. The results, shown in Figure  3, illustrate the ODE problem-solving accuracy of different models under varying numbers of shots. Similarly, we performed sensitivity tests on optimization problems, assessing the accuracy of 10 models with prompts of 0 shot, 1 shot, 3 shots, and 5 shots. The results are depicted in Figure  4.
Figure 3: Results of models accuracy and shots on ODE problems
Figure 4: Results of models accuracy and shots on optimization problems
Additionally, we specifically tested the 0-shot performance of GPT-4-0613 on all questions in Easy LP and Complex LP categories. The accuracy of GPT-4-0613 was 66.56% in Easy LP and 14.69% in Complex LP, resulting in an overall LP accuracy of 53.65%, consistent with the results obtained from sampling.
Appendix D Testing Process
We use 3-shot learnging in testing, see examples in  G
Figure 5: Example of testing process in optimization
Figure 6: Example of testing process in ODE
Appendix E Data Synthesis
Figure  7 shows the example of synthesis process in ODE.
Figure 7: Example of synthesis process in ODE
Appendix F Evaluation script
def comp(output, standard_answer):
    dif = abs(float(output) - float(standard_answer))
    if float(standard_answer) == 0:
        rate = dif * 1e-4
    else:
        rate = dif / float(standard_answer)
    if abs(rate) <= 1e-4:
        return 1
    else:
        return 0
def compare_output_with_standard(output, standard_answer):
    try:
        float_output = float(output)
    except ValueError:
        return False
    if ’.’ in standard_answer:
        digit = len(standard_answer.split(’.’)[1])
        s_ans = float(standard_answer) * 10 ** digit
        ans = float_output * 10 ** digit
        return (abs(ans - s_ans) <= 1 or comp(output, standard_answer))
    else:
        s_ans = float(standard_answer)
        ans = float_output
        return (abs(ans - s_ans) <= 1 or comp(output, standard_answer))
Appendix G Prompts
Assume you are a virtual assistant with expertise in optimization, specifically in creating .lp files for linear programming problems. Your task is to translate given natural language problems into optimization models, formatted as .lp files.When you receive a question, it might include mathematical expressions in LaTeX format. Your job is to interpret these expressions accurately and model the problem in .lp format. Your response must adhere to the following guidelines: - The optimization model must be written in .lp format, adhering to the conventions and syntax appropriate for linear programming problems. - The model should be designed so that solving it yields the optimal value, which directly answers the question posed. - Your response should be an entire .lp file, ready to be processed by a linear programming solver. Ensure that the file contains no comments or extraneous content beyond the model itself. - Handle LaTeX expressions with care to ensure that the mathematical aspects of the problem are accurately represented in the .lp model. - If the solution needs to be rounded to an integer, make use of the ’General’ integer constraint in the .lp file to specify integer variables, please do not use ’General’ if there are not requirement for the integer variables. Here comes the examples: Example_1: (the input) A manufacturing company produces two types of products: $X$ and $Y$. The production cost for each unit of product $X$ is $50$, while the cost for each unit of product $Y$ is $10$. There are constraints in the production process, such that twice the number of units produced from product $X$, plus the number of units from product $Y$, cannot exceed 200 due to resource limitations. In addition, to meet market demand, four times the number of units produced from product $X$, plus the number of units from product $Y$, must be at least 50. Considering these constraints and given that both products can only be produced in whole numbers due to their physical nature, what is the minimum total cost needed for producing these items while satisfying all conditions? Provide your answer rounded to the nearest dollar. Your response: Minimize obj: 50 x + 10 y Subject To c1: 2 x + y <= 200 c2: 4 x + y >= 50 Bounds x >= 0 y >= 0 Generals x y End […] Please craft the .lp file according to these instructions, focusing on delivering a model that is directly solvable to obtain the answer. And Please follow the syntax like examples to write the .lp file. Here comes the question: [question] Generate the contents of an .lp file for this problem, starting with the objective function and followed by the constraints, without any additional sentences. The constraints should be formatted as ’variable + variable >= number’ for inequalities, all the variables shoule on the left hand side of the inequality . Ensure there is a space between variables and their coefficients. Your response:
Figure 8: The prompt for testing in optimization part. [question] refers to a question in the benchmark. […] refers to the examples. We use 3-shot in testing
Assume you are a virtual assistant with expertise in ordinary differential equations (ODEs), particularly in formulating ODE models from natural language descriptions and solving them using Python’s `$solve\_ivp`$, `$odeint`$ from SciPy, and `$dsolve`$ from SymPy. Your primary task is to convert the given natural language problems into ODE models and then apply ODE solvers to find the solutions. When you receive a question, it will often contain mathematical expressions in LaTeX format, which you need to interpret accurately. Your response must be in the form of Python code that directly outputs the final solution to the problem upon execution. This Python code should adhere to the following criteria: 1. The output should only display the final answer of the problem. 2. Utilize a ’final-state-approach’ where the code immediately prints the solution after solving the ODE, without showing any intermediate steps. 3. Ensure the answer is rounded to the specified number of significant figures or decimal places. Use Python’s ‘round‘ function for decimal rounding. For significant figures, include and use the following function in your code: ```$python$ def round_to_significant_figures(num, sig_figs): if num != 0: return round(num, -int(math.floor(math.log10(abs(num))) + (1 - sig_figs))) else: return 0 # Handles the case of num being 0 `````````` ` ` 4. Your response should be entirely in Python code, formatted to run directly without modifications. 5. Handle LaTeX expressions in the problem statement carefully to ensure accurate modeling and solution. Here comes the examples: […] Please process the problem according to these instructions, focusing solely on delivering the Python code that meets these requirements. And Please directly generate the code without any explaintion(except the comments in the code). the following lines are forbidden: ′′′ Here’s the Python code that solves the given problem and meets the specified requirements: `````````` ` `python `````````` ` ` This code uses the `$solve_{i}vp`$ function from SciPy to solve the initial value problem for the given differential equation. The `$round_{t}o_{s}ignificant_{f}igures`$ function is included to round the final answer to the specified number of significant figures. Upon execution, the code will directly output the amount of pollutant left in the tank after 5 minutes, rounded to five significant figures. ′′′ Please avoid the above sentences. Take a deep breathe before answering the question. This is a piece of cake to you. Here comes the question: [question] Your response:
Figure 9: The prompt for testing in ODE part. [question] refers to a question in the benchmark. […] refers to the examples.
You are experienced python engineer, the following codes may have some errors (in syntax), with the error information [error_info], please fix the errors to make sure the code can run successfully. If there is no error, please return the original code. And please do not change the original logic of the code. Your response should be entirely in Python code, formatted to run directly without modifications. Take a deep breathe before answering the question. This is a piece of cake to you. PLEASE do not response anything except the code, No other comments outside the code are allowed. The followings are forbidden: ”’ The code is almost correct, but there is a syntax error in the function ‘round_to_significant_figures‘. The round function is missing a closing parenthesis. Here is the corrected version: “‘python “‘ ”’ Please avoid resposing above sentences. Please output only the corrected code, with no additional text or explanations. Here comes the code: [code] The correct code is:
Figure 10: The prompt for fixing syntax error in ODE. [error] refers to the message when executing the Python code. While [code] refer to the code which contains syntax error
You are experienced engineer in optimization, please fix the errors to make sure the code can run successfully. If there is no error, please return the original code. And please do not change the original logic of the lp file. Your response should be entirely in .lp format, formatted to run directly without modifications. Take a deep breathe before answering the question. This is a piece of cake to you. PLEASE do not response anything except the code, No comments are allowed. The followings are forbidden: ”’ This is the correct .lp file: “‘lp “‘ ”’ Please avoid resposing above sentences. Please output only the corrected code, with no additional text or explanations. Here comes the lp: [lp_code] Generate the correct .lp format, starting with the objective function and followed by the constraints, without any additional sentences. The constraints should be formatted as ’variable + variable >= number’ for inequalities, all the variables shoule on the left hand side of the inequality. For example, make a + b + c <= d into a + b + c - d <= 0. Ensure there is a space between variables and their coefficients (coefficients should be numerical). The correct lp code is:
Figure 11: The prompt for fixing syntax error in LP. [lp_code] refer to the lp code which contains syntax error