### v2

# Mamo: a Mathematical Modeling Benchmark with Solvers

Abstract
Mathematical modeling involves representing real-world phenomena, systems, or problems using mathematical expressions and equations to analyze, understand, and predict their behavior. Given that this process typically requires experienced experts, there is an interest in exploring whether Large Language Models (LLMs) can undertake mathematical modeling to potentially decrease human labor. To evaluate of LLMs in mathematical modeling, we introduce a new benchmark, Mamo, that transcends traditional result-oriented assessments. Unlike conventional methods that primarily assess LLMs based on the accuracy of solutions to mathematical problems, our approach offers deeper insight into the modeling process itself. By focusing on the processes LLMs undertake rather than the correctness of their final solutions, Mamo pioneers a novel evaluation paradigm. This shift underscores the importance of understanding the inherent modeling capabilities of LLMs, paving the way for a more nuanced and comprehensive analysis of their problem-solving strategies. Our work marks a significant advancement in the field, suggesting a new direction for future research by emphasizing the evaluation of LLMs’ modeling processes over the mere correctness of answers. This benchmark not only facilitates a better understanding of LLMs’ mathematical modeling capabilities but also sets a new standard for evaluating their performance in complex problem-solving scenarios.
1 Introduction
Recent advancements in Large Language Models have garnered widespread interest, demonstrating remarkable capabilities across a broad spectrum of natural language processing tasks Ouyang et al. (2022); OpenAI (2023); Nijkamp et al. (2022); Tang et al. (2024). However, the domain of mathematics remains a critical aspect of LLM evaluation Wei et al. (2022). This focus not only gauges LLMs’ proficiency in comprehending and addressing mathematical challenges but also serves as a profound indicator of their underlying abilities in abstract conceptualization and logical reasoning, among other intellectual faculties.
Natural language captures human complexity and nuance, while mathematical language excels in precision and modeling the natural world (Giordano et al., 2013). Bridging their gap requires high-level cognitive skills akin to Artificial General Intelligence (AGI), a task at which LLMs show promise. However, the challenge lies in the evaluation of modeling, as perfectly representing reality may be inherently elusive.
In this endeavor, solvers [^1] In optimization, solvers are algorithms used to find the best solution to a problem. They search through possible options to maximize or minimize an objective under given constraints if any. By using solvers, The challenge becomes problem formulation under a given solver or a solver set. becomes critically important as it standardizes the modeling output and could provide a way to falsifiably validate the produced answers. Recent studies have already combined LLMs with solvers Yang et al. (2023); Feng et al. (2023); Pan et al. (2023). In the field of optimization, AhmadiTeshnizi et al. (2023b) introduced OptiMUS, an LLM-based agent that uses a solver to address optimization problems. The key task for LLM in OptiMUS involves first generating a mathematical formulation of a real-world problem, followed by writing Python code to invoke a solver. This implies the potential for testing LLM’s modeling capabilities.
We thus develope Mamo, a new benchmark that leverages solvers to assess the mathematical modeling prowess of LLMs. Our approach shifts from conventional outcome-focused evaluation to process-focused evaluation, offering in-depth insights into the LLMs’ problem-solving strategies. We concentrate on the models’ mathematical modeling skills, delegating the task of solving abstract problems to solvers to circumvent errors arising from the models’ computational capabilities.
The contributions of this paper are as below: (1) We introduce a new benchmarking strategy to assess mathematical modeling through Solvers, facilitating a rigorous evaluation process. (2) We have developed a new benchmark, named Mamo, specifically tailored for the evaluation of mathematical modeling capabilities. Mamo encompasses a wide array of modeling questions (With a total of 1209 meticulously curated questions), including ordinary differential equations and optimization problems within linear programming and mixed-integer linear programming frameworks.
2 Related Work
Recent studies have made significant strides in applying LLMs to mathematical problem-solving, introducing innovative datasets for comprehensive evaluation. Frieder et al. (2024) and Yuan et al. (2023) have focused on datasets that test LLMs across various math problems and arithmetic expressions, respectively. Lewkowycz et al. (2022) explored training models using natural language paired with LaTeX-formatted mathematical content from arXiv, enhancing syntactic understanding. Zhang et al. (2024) developed the CARP dataset for computation-intensive challenges, while He et al. (2024) introduced OlympiadBench, a bilingual and multimodal benchmark for competition-level mathematical reasoning and proofs. Liu et al. (2024) constructed MathBench, which is designed with a carefully structured difficulty hierarchy and focuses on assessing the understanding of theoretical knowledge in LLMs. AhmadiTeshnizi et al. (2023a) constructed NLP4LP, a benchmark comprising 52 linear programming (LP) and mixed-integer linear programming (MILP) problems. These efforts collectively push the boundaries of LLMs in mathematical cognition, showcasing the growing complexity and specificity of tasks LLMs are being trained to tackle.
The pursuit of enhancing and devising novel solvers has seen a variety of research efforts, with a notable focus on the integration of LLMs. Yang et al. (2023) investigated the potential of employing LLMs as direct solvers, pioneering the exploration of LLMs’ autonomous problem-solving capabilities. He-Yueya et al. (2023) introduced a hybrid approach that marries an LLM with an external symbolic solver, specifically for equation solving, marking a significant advancement in combining neural and symbolic computation. Furthermore, Pan et al. (2023) developed LOGIC-LM, a cutting-edge framework that fuses LLMs with symbolic solvers, aimed at enhancing the solving of logical problems. In a closely related study, AhmadiTeshnizi et al. (2023a) introduced OptiMUS, an LLM-based agent explicitly crafted for formulating and solving optimization problems, thus demonstrating the broad applicability and potential of LLMs in tackling intricate computational challenges. These endeavors underscore the innovative directions being explored in the realm of solver development, leveraging the robust capabilities of LLMs.
3 Background
3.1 Mathematical Modeling
Giordano et al. (2013) emphasize the role of mathematical models as intermediaries that transform real-world problems into structured mathematical forms, facilitating the discovery of solutions. Despite LLMs’ proven competencies across various tasks, their capacity for developing mathematical models remains uncertain. This involves two key phases: model formulation, which demands a thorough comprehension of the problem, and model resolution, typically addressed by computational solvers. To accurately gauge LLMs’ potential in this domain, our strategy focuses on their ability to construct mathematical models, entrusting calculations to specialized solvers. This approach highlights LLMs’ creativity and generalization skills, simultaneously capitalizing on solvers’ computational efficiency.
Motivationsto Benchmark Mathematical Modeling for LLMs   We evaluate mathematical models for two reasons. The first reason is out of curiosity: We aim to investigate whether LLMs can do what expert humans can. Mathematical modeling is considered a highly advanced capability that we believe experts possess. This is a beneficial test for exploring the boundaries of LLMs’ capabilities and to what extent they can approach human intelligence. On the other hand, in terms of evaluating the models themselves, most current benchmarks, especially those related to mathematics like GSM8K Cobbe et al. (2021), etc., show that the performance has become saturated. Even a 7B model Yu et al. (2023b); Liu et al. (2023) (sometimes with a verifier-like model Yu et al. (2023a)) can achieve a accuracy exceeding 80%, closely approaching that of GPT 4. Therefore, we believe there is a need to construct more challenging mathematical datasets. Traditional mathematical modeling is a labor-intensive, custom process reliant on human expertise, raising the question: Can large language models automate or assist in this process?
3.2 Using Solvers in Mathematical Modeling
The model generated by the LLM is then subjected to verification against real-world data to assess its validity, ensuring it addresses the problem logically and sensibly. The solver, an algorithmic tool designed to find solutions to mathematical models, is also involved in this process and it is also beneficial for verification.
Different types of solvers are specialized for various classes of problems, such as optimization solvers. However, they share the common property: They are all algorithmic tools designed to find solutions to mathematical models.
Rationale of Using solver in Mathematical Modeling for LLMs These solvers operationalize the abstract constructs created by LLMs, translating them into tangible outputs that can be analyzed and validated against real-world data and phenomena. The solver is a crucial instrument that resolves the model presented by the LLM, and by comparing the solution it provides with the correct answer, we assess the accuracy of the mathematical model generated by the LLM.
4 Philosophy of Mamo Benchmark
4.1 Methodology: exact answer verification with solvers
We utilize solver to resolve this mathematical model. The solver operates at the final state of the question-solving process, meaning it directly computes the final answer from the model without intermediate steps. By executing the solver and juxtaposing its output ($\widehat{A}$) with the correct answer (A) to the original problem, we can verify the accuracy of the Language Model’s mathematical model(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$), since typically the wrong ${\widehat{M}}$ will result in wrong ($\widehat{A}$), like the examples in Appendix  G. This final-state approach ensures that our assessment is squarely focused on the Language Model’s capacity to model the problem, abstracting from any additional computational steps that might otherwise obscure the evaluation of the model’s validity.
Figure 1: The pipeline to use exact answer verification via an additional solver.
4.2 Scope of Mathematical Modeling
We construct the category of Mathematical Modeling by referring to Giordano et al. (2013). In the context of Natural Language we propose a framework in mathematical modelling in LLM, specified in solving natural language problem, see Table 7 in Appendix F.
Given the broad spectrum of mathematical modeling and the availability of specific solvers, our benchmark is meticulously designed to encompass domains where these tools can be effectively utilized. This section outlines the rationale for our focus on certain areas of mathematical modeling, based on solver availability and the clarity of the modeling process.
Availability of Sophisticated Solvers: Advanced optimization solvers, such as COPT Ge et al. (2023), along with Python libraries for solving ordinary differential equations (ODEs), offer a strong foundation for testing the capabilities of Large Language Models (LLMs). These tools enable a detailed and automated evaluation of the LLMs’ abilities to abstract, formulate, and accurately solve mathematical problems within these specific domains.
Challenges in Modeling Changes and Differences, Proportionality and Geometric Similarity: Although simpler tools like calculators could address problems related to modeling changes and differences, proportionality and geometric similarity, the lack of a defined, structured modeling process complicates the implementation of our final-state-approach benchmark. Our objective is to assess the LLMs’ skill in developing comprehensive mathematical models that adhere to a solvable framework. The absence of a distinct modeling procedure in these areas impedes a standardized and objective assessment of an LLM’s modeling efficiency.
Following the considerations mentioned above, we have developed our benchmark in the fields of optimization, and ODE, due to the lack of advanced and universal Partial Differential Equations (PDE) solvers. In the optimization segment, our focus is on Linear Programming (LP) and Mixed-Integer Linear Programming (MILP) problems. Compared to NLP4LP, introduced by AhmadiTeshnizi et al. (2023b) and focusing exclusively on optimization, our benchmark includes a larger number of problems. This benchmark set is highly scalable, with potential future expansions into nonlinear optimization and probabilistic modeling. Additionally, as new and advanced solvers emerge, it could be expanded into areas like Partial Differential Equations (PDE) and beyond.
5 Data Collection
5.1 Data Selection and Source Credibility
The dataset for our benchmark consists of a blend of manually selected and GPT-synthetized questions. All items within the data set have been subjected to a meticulous review process conducted by qualified collectors and domain experts to ensure their validity and relevance. The data collection and synthesis process is conducted by a team of qualified individuals with extensive expertise in various mathematical disciplines  [^2] See the details of the annotator qualifications in Appendix C.
Manually Selected Questions
The manually selected questions are high quality and reliable. (1)Source and Reliability: When doing manually selecting, we refer to the textbooks, exercises and the solutions in ODE and optimization area(Thomas et al., 2016; Stewart et al., 2014; Lial et al., 2017; Braun, 1993; Boyce & DiPrima, 2012; Giordano et al., 2014; Zill, 2013; Giordano et al., 2014; Bertsimas & Tsitsiklis, 1997; Hurlbert, 2010). The selecting process focus on deriving the mathematical model in the questions and assign a reasonable scenario. The source ensure the correctness of the answer of each question. (2) Transformation: Each question is meticulously transformed to fit the benchmark’s criteria(the properties in Section 5.3), which may involve assigning new scenarios or adapting the question and its answer to align with our testing methodology.
Question Formulation and Verification
The process of synthesizing data starts with the creation of typical mathematical constructs, into which random parameters are introduced to give them specificity. This is followed by the computation of answers to establish a ground truth. Subsequently, GPT-4 is employed to craft real-life scenarios based on these predefined models, effectively translating abstract mathematical concepts into tangible, context-rich problems. This approach not only tests the model’s ability to handle mathematical reasoning but also its capacity to contextualize mathematical principles within real-world situations. See the example in Appendix K.
5.2 Data Quality Checking
In the ODE part, we employed GPT-4 to confirm the solvability of newly generated questions, filtering out invalid ones. A thorough evaluation identified questions lacking information, with incorrect answers, and unclear statements. Consequently, problematic questions were deleted or corrected, resulting in a refined dataset. In the optimization section, particularly with the easy LP part, invalid questions were filtered out, and those with incorrect answers or misleading descriptions were corrected. In the complex LP part, questions formulated with wrong mathematical models were corrected. The summary of the review process is presented in Table 1.
Category Initial Questions Filtered Final Questions
ODE Part 383 37 346
Easy LP Part 688 48 640
Complex LP Part 211 0 211
Table 1: Summary of the review process.
Reviewer Accuracy Rate
Reviewer 1 90.0%
Reviewer 2 84.0%
Reviewer 3 88.0%
Reviewer 4 74.0%
Table 2: Accuracy of reviewers.
For the cross-review process, 50 questions from the ODE dataset were selected for deeper analysis by four independent reviewers. Reviewers evaluated the questions, marking responses as either numeric answers or "error" if the question was problematic. The effectiveness of this review was quantified using metrics like Cohen’s Kappa to measure inter-rater reliability and accuracy rates for individual reviewers. The average Cohen’s Kappa was 0.60, indicating moderate to substantial agreement among reviewers, with individual Kappa values ranging from 0.50 to 0.71. Accuracy rates are moderately high as shown in Tables 2. These results highlight both the effectiveness of the review process and areas where alignment and training can improve consistency and accuracy among reviewers. See the review details in Appendix B.
5.3 Guidance of Data Collection
The collection of problems for our benchmark are governed by specific criteria designed to evaluate the mathematical modeling capacity of Large Language Models (LLMs). The guidelines ensure that problems are appropriate for our final-state approach and yield numerical answers suitable for automated verification. The following are the key considerations in our benchmark guidance:
Property 1.
Final-State Approach Solvability  Problems should be structured so that they can be solved by a solver in a final-state approach, meaning the solution process should not require intermediate steps once the mathematical model has been formulated by the LLM.
Property 2.
Unified and Numerical Answers: To handle potentially various valid solutions for one problem: For Ordinary Differential Equations (ODEs), questions should focus on determining the value of the original function at a specific time $t_{0}$. For optimization problems, questions should aim to find the optimal value. These constraints facilitate the testing accuracy and the convenience of verifying potentially multiple valid solutions to a single problem. Besides, the answers to all questions must be numerical, allowing for clear-cut evaluation and comparison with the solver’s output.
Property 3.
Significant Figures and Precision  Each question will explicitly state the required significant figures or the level of numerical precision for the answer. This clarity ensures that the LLM’s output can be properly evaluated against the solver’s results, maintaining consistency and rigor in the assessment of the model’s accuracy.
To assess the LLM’s capability to abstract mathematical models from real-world scenarios, we introduce Property 4.
Property 4.
Real-World Problem Context  Questions should be framed as real-world problems without explicitly stating the underlying mathematical model, challenging the LLM to demonstrate its capacity for abstraction and application in practical scenarios.
5.4 Data Statistics
In the section on Ordinary Differential Equations (ODE), we present a total of 346 problems: 196 are based on first-order equations, 110 on second-order equations, and 40 on systems of equations, offering a comprehensive exploration across different levels. The optimization section is divided into two segments: Easy LP, featuring 652 high school-level Mixed Integer Linear Programming (MILP) problems, and Complex LP, comprising 211 undergraduate-level problems that integrate both LP and MILP. See the word cloud of the combined data in the Appendix L.
6 Evaluation
6.1 Evaluation Protocol
6.1.1 Evaluation Details
Our evaluation protocol is meticulously designed to test the LLMs’ mathematical modeling process. For ODE problems, the LLM is prompted to generate Python code from a natural language description. The code is executed, and its output is compared with the correct answer to assess accuracy. In optimization problems, the LLM is asked to express the model in .lp format, after which the COPT solver is used to find the optimal value for comparison with the correct answer.
Model output
In ODE parts, we ask LLMs to output Python code calling solvers ( solve_ivp odeint in SciPy, and dsolve in NumPy). In the optimization sections, we prompt LLMs to generate the standard .lp file. Compared to the testing methodology in NLP4LP (AhmadiTeshnizi et al., 2023b), which involves prompting the model to output Python code for invoking the solver, the .lp format more closely aligns with the standard optimization form, making it more readable. Furthermore, optimization solvers such as COPT (Ge et al., 2023) can directly read and solve the .lp file, minimizing potential issues related to the LLM’s coding capabilities. To better adhere to the format, we conducted tests using 3-shot learning. Detailed information about the number of few-shot samples used in our experiments and the result can be found in the Appendix (see Appendix I and Section O). Besides, at the beginning of the testing, we do the first clean of the codes, clearing up the string like ‘ “‘ python “‘ ’ that will affect our testing.
Mitigation of Formatting Errors
It is essential to distinguish errors in modeling from those in coding or file formatting [^3] Appendix  H displays the example of formating errors instead of modeling errors.. Therefore, if the LLM’s output contains syntax errors in python code or .lp code, we need to rectify these issues without altering the underlying logic of the code. This step ensures that our evaluation focuses solely on the modeling process. To achieve this, we use a language model to fix any syntax error, referring to it as the code modifier. In our experiments, we use the original language model and GPT-4-0613 as code modifiers. For the sake of reproducibility, we choose the original LLM; To ensure the comparison focuses exclusively on modeling capacity, we choose GPT-4-0613 as the code modifier due to its exceptional coding proficiency and operational reliability, as well as its stability than other provisional release Language Models. See the comparison of the result from different code modifiers in Section 6.3, and prompts in Appendix  O.
6.1.2 Evaluation Metrics
Given the potential for minor discrepancies in numerical solutions, our protocol includes a comparison rule that accommodates for slight inaccuracies. Let $O$ denote the output from the LLM, and $A$ denote the standard answer. The comparison is defined by the following criteria  444 See the evaluation script in Appendix M
General Comparison Criterion Let $n$ denotes the decimal places of $A$ ($n=0$ if $A$ is integer). The Answer $O$ is correct if
$\left|\frac{O-A}{A}\right|<10^{-4}\text{ or }|O-A|<\min\{10^{-n},10^{-2}\}$ (1)
This metric is designed [^5] See the case study in Appendix D to account for the precision errors inherent in solvers and the specified significant figures in questions, ensuring that the LLM’s output is accurately evaluated against the standard answer with allowances for numerical precision and variability in computational responses.
6.2 Benchmarking Results
Table 3: Evaluation Results on the Mamo Benchmark. The subscript denotes improvement using the original LLM as a code modifier, which corrects syntax errors without changing the underlying logic, differentiating modeling from formatting errors (see Section 6.1.1). The "Overall" score represents the weighted average of correct rates, weighted by question count. The highest score in each category is marked in boldface; an underline signifies the top score with the LLM itself as code modifier. The score after using LLM as code modifier is computed as the sum of the original score and the value in the subscript.
Models ODE Optimization Overall (%)
First Second System Easy Complex
order (%) order (%) (%) LP (%) LP (%)
GPT-4o 67.86 +0.51 36.36 +0.91 40.00
+0.00
87.27
+0.15
22.75
+8.53
66.67 +1.73
GPT-4-turbo 64.80
+1.02
32.73
+0.91
40.00
+0.00
87.88 +0.00 23.22
+6.64
66.34
+1.41
GPT-4 59.18
+1.02
28.18
+0.91
40.00
+2.50
86.50
+0.46
21.08
+3.09
63.81
+1.12
GPT-3.5-turbo 16.33
+6.12
7.27
+4.55
10.00
+5.00
81.29
+3.53
9.48
+0.00
49.13
+3.48
Claude-3-Opus 55.61
+3.57
33.64
+1.81
45.00
+2.50
83.74
+0.31
9.48
+0.47
60.38
+1.08
Claude-3-Sonnet 52.04
+2.04
24.55
+0.90
27.50
+5.00
83.59
+0.61
24.64
+0.48
60.96
+0.99
Claude-3-Haiku 41.33
+9.18
23.64
+0.91
17.50
+2.50
78.53
+7.36
17.54
+1.42
54.84
+5.87
Gemini-1-pro 25.00
+8.16
13.64
+0.00
2.50
+2.50
78.83
+1.08
14.69
+0.00
50.45
+1.99
Gemini-1.5-pro 64.80
+1.02
30.00
+0.00
45.00 +5.00 85.28
+0.00
30.81 +5.21 66.09
+1.24
DeepSeek-v2 39.29
+1.02
20.00
+9.09
30.00
+0.00
86.50
+0.16
3.79
+0.00
56.49
+1.08
DeepSeek-coder-33B 28.57
+2.04
11.82
+0.00
12.50
+0.00
75.77
+1.22
18.96
+0.00
50.29
+0.99
DeepSeek-math-7b-base 3.57
+0.00
2.73
+0.00
2.50
+0.00
49.39
+1.22
6.64
+0.00
28.70
+0.66
DeepSeek-math-7b-rl 1.02
+0.51
1.82
+0.00
5.00
+0.00
20.55
+0.62
0.00
+0.47
11.58
+0.50
Llama-3-8b 0.51
+0.51
0.00
+0.00
0.00
+0.00
55.37
+0.46
6.64
+0.00
31.10
+0.33
Llama-3-70b 4.08
+0.51
1.82
+1.82
2.50
+0.00
66.10
+0.62
9.00
+0.00
38.13
+0.58
Qwen-1.5-72b-Chat 5.61
+6.12
4.55
+0.00
5.00
+7.50
65.03
+6.29
5.69
+0.00
37.55
+4.64
Qwen-1.5-110b-Chat 28.57
+4.08
6.36
+1.82
7.50
+0.00
65.18
+10.89
13.27
+2.37
43.93
+6.11
Mixtral-8x7b-instruct 23.47
+4.59
4.55
+1.81
10.00
+0.00
67.79
+2.15
9.95
+0.00
42.85
+2.06
Mixtral-8x22b-instruct 52.04
+1.53
11.82
+1.82
20.00
+0.00
81.90
+2.61
19.91
+1.42
57.82
+2.06
  • 1
    Here GPT-4-turbo denotes GPT-4-turbo-2024-04-09; GPT-4o denotes GPT-4o-2024-05-13; GPT-4 refers to GPT-4-0613; GPT-3.5-turbo denotes GPT-3.5-turbo-0125
Our evaluation  [^6] See the setting in Appendix N includes both closed-source LLMs and open-source LLMs. The closed-source commercial LLMs include the GPT-4 series OpenAI (2023), the Claude series  [^7] https://www.anthropic.com/news/claude-3-family, and the Gemini series  [^8] https://deepmind.google/technologies/gemini/. The open-source LLMs include DeepSeek series by  DeepSeek-AI (2024),  Guo et al. (2024) and  Shao et al. (2024); Llama-3 models  [^9] https://github.com/meta-llama/llama3; Qwen-1.5 (Bai et al.) and Mixtral (Jiang et al.). Table 3 presents the evaluation results of different language models on the Mamo Benchmark. The performance of these models is assessed both in their original form and when enhanced with the language model itself to rectify syntax errors (the improvement is indicated by the subscript). Table 5 in Appendix A shows results when GPT-4-0613 is selected as the code modifier.
The evaluation results on the Mamo Benchmark reveal that GPT-4o stands out among the models in both ODE and LP tasks. Specifically, GPT-4o achieves the highest performance in ODE tasks with 68.37% in first-order and 37.27% in second-order equations, and also excels in LP tasks with 87.42% in easy LP and 31.28% in complex LP scenarios when using itself as code modifier. Additionally, DeepSeek-v2, as an open-source LLM, demonstrates significant capability, achieving a 57.57% success rate across all tasks when enhanced with syntax correction by itself, underscoring its potential despite its open-source nature. Moreover, Mixtral-8x22B also exhibits competitive performance relative to DeepSeek-v2. After fix syntax error by itself, it achieves a remarkable score of 21.33% on the complex LP task, showcasing its potential in modeling complex optimization problems.
The inclusion of the LLM itself as code modifier to rectify syntax errors results in significant improvements across various models. For instance, GPT-3.5-turbo’s performance in ODE tasks increased from 12.72% to 18.21% overall and Claude-3-Haiku shows a remarkable improvement in ODE tasks from 32.95% to 38.73%. Similarly, DeepSeek-v2, as an open-source LLM, demonstrates significant capability with its overall ODE performance increasing from 32.08% to 35.55% when enhanced with syntax correction.
6.3 On the Format Errors and Code Modifiers
Table 4: Execution Success Rates for Raw, Self-Modified, and GPT-4 Modified Methods, denoting scenarios without code modifiers, utilizing the LLM as a code modifier, and employing GPT-4-0613 as the code modifier. This table highlights the success rates of file execution (correct format) across these three scenarios, differing from the Table 3 which shows the correctness rate.
Models ODE (%) Optimization (%)
raw self-modified GPT-4-modified raw self-modified GPT-4-modified
GPT-4o 96.82 99.42 99.42 92.93 97.91 97.91
GPT-4-turbo 96.82 99.13 98.84 93.86 97.80 97.56
GPT-4 94.51 97.69 97.69 90.96 95.13 95.13
GPT-3.5-turbo 37.28 65.03 93.93 87.95 89.80 96.76
Claude-3-Sonnet 84.68 93.06 97.40 91.89 93.97 96.76
Claude-3-Haiku 78.03 92.77 98.27 82.61 92.58 96.29
Claude-3-Opus 88.73 96.82 97.98 90.38 93.40 95.94
Gemini-1.0-pro 56.36 69.94 88.73 88.06 89.11 94.21
Gemini-1.5-pro 93.64 98.27 98.84 95.48 98.73 96.87
DeepSeek-v2 70.81 83.82 96.24 90.27 92.00 94.90
DeepSeek-coder 88.44 93.64 95.95 92.58 96.29 98.84
DeepSeek-math-7B-base 28.90 29.77 79.77 55.50 58.40 75.67
DeepSeek-math-7B-rl 5.20 6.64 80.64 19.93 21.55 55.97
Llama-3-8B 20.23 26.01 68.21 75.09 76.36 79.14
Llama-3-70B 24.56 30.92 77.46 77.64 79.49 89.46
Qwen-1.5-72B-Chat 55.49 89.60 95.38 77.40 92.24 95.37
Qwen-1.5-110B-Chat 51.45 63.29 95.09 71.26 84.01 90.15
Mixtral-8x7B-instruct 50.87 69.08 91.62 84.13 87.83 96.29
Mixtral-8x22B-instruct 80.92 88.73 96.24 87.72 93.51 96.29
The data presented in Table 4 provides a comparative analysis of the performance (rate of correct format) of various models using different modification methods. The table considers three scenarios: no modification, self-modification, and modification provided by GPT-4-0613.
The GPT-3.5-turbo model shows the most dramatic improvement when modified by GPT-4, jumping from a 37.28% success rate to 93.93%. Interestingly, Claude-3 models (Sonnet and Haiku) and DeepSeek models (DeepSeek-v2, DeepSeek-coder) also exhibit high improvement rates when enhanced by GPT-4-0613, highlighting the compatibility of GPT-4’s modification methods with other models. In the realm of Optimization tasks, the GPT-3.5-turbo model again shows a significant improvement when enhanced by GPT-4, jumping from a 87.95% success rate to 96.76%.
The modification from GPT-4 is generally superior to both self-modification and no modification. This indicates that GPT-4 can serve as a more effective and fair code modifier. By using GPT-4 for syntax error correction, we can better examine the modeling ability of different LLMs without the confounding factor of formatting errors. Table 5 in Appendix A shows the results of selecting GPT-4-0613 as code modifier.
6.4 Analysis
Given the results in Section 6.26.3 and Appendix A, several observations emerge from the benchmark results. For open-source models, larger models generally perform better, as evidenced by comparisons between Llama-3-70B and Llama-3-8B, Mixtral-8x22B-instruct and Mixtral-8x7B-instruct, with the larger models consistently achieving higher scores.
From Table 4, it is apparent that some models struggle with adhering to the syntax of the solver, such as GPT-3.5-turbo and certain open-source models. Some models like Llama-3-8B may not closely follow the instructions, even worse than the syntax error.These models tend to show significant improvements in both correct rates (as indicated by the subscript in Table 3) and execution success rates after the application of code modifiers, particularly when their initial performance without modifiers is low. Notably, in the ODE part, GPT-3.5-turbo achieves a remarkable 28.03% increase in execution success rate by self-modified, accompanied by substantial improvements in ODE performance. In Optimization part, Qwen-1.5-72B-Chat follows the same trends. When comparing the results from Table 3 and Table 5, it becomes clear that the improvements are even more pronounced for open-source models when utilizing GPT-4 modifications rather than self-modifications. See Appendix E for the analysis of typical modeling errors and case study,
7 Conclusion and Future Directions
In conclusion, Mamo’s innovative benchmark shifts the evaluation focus of LLMs towards the modeling process, offering a nuanced assessment of their mathematical problem-solving capabilities. By isolating modeling from solving through the use of solvers, Mamo provides a comprehensive understanding of LLMs’ strengths and limitations in mathematical modeling. This approach not only enhances our insights into the fundamental modeling abilities of LLMs but also guides future research by highlighting the importance of the modeling process in solving mathematical problems.
Appendix A GPT-4-0613 as Code Modifier
[…]
Models ODE Optimization Overall (%)
First Second System Easy Complex
order (%) order (%) (%) LP (%) LP (%)
GPT-4o †† 68.37 37.27 40.00 87.42 29.38 68.07
GPT-4-turbo †† 65.82 33.64 40.00 88.19 27.96 67.58
GPT-4 †† 60.20 29.09 42.50 86.96 24.17 64.93
GPT-3.5-turbo †† 44.90 20.00 20.00 85.43 9.48 57.48
Claude-3-Opus †† 59.18 36.36 47.50 84.20 14.22 62.37
Claude-3-Sonnet †† 56.63 26.36 32.50 85.12 27.96 63.44
Claude-3-Haiku †† 51.53 25.45 22.50 86.81 18.96 61.54
Gemini-1-pro †† 40.31 18.18 10.00 80.21 15.64 54.51
Gemini-1.5-pro †† 65.82 30.91 50.00 85.28 34.12 67.08
DeepSeek-v2 †† 54.08 30.91 30.00 86.81 3.79 60.05
DeepSeek-coder †† 30.10 11.82 15.00 78.83 18.96 52.27
DeepSeek-math-7b-base †† 13.78 9.09 5.00 59.05 9.48 36.72
DeepSeek-math-7b-rl †† 13.78 7.27 7.50 47.70 1.90 29.20
Llama-3-8b †† 11.22 0.91 0.00 63.19 8.53 37.47
Llama-3-70b †† 14.80 7.27 10.00 70.71 9.48 43.18
Qwen-1.5-72b-Chat †† 16.84 13.01 5.45 78.53 6.64 47.80
Qwen-1.5-110b-Chat †† 40.82 13.64 20.00 78.83 14.22 53.51
Mixtral-8x7b-instruct †† 32.14 11.82 20.00 75.46 9.95 49.38
Mixtral-8x22b-instruct †† 55.10 17.27 22.50 85.43 22.27 61.21
Selecting GPT-4 as a code modifier generally results in better performance than using the LLM itself for self-improvement. For instance, DeepSeek-v2’s accuracy in ODE tasks improves from 32.08% with self-modification to 43.93% when GPT-4 is used as the code modifier. This pattern indicates that GPT-4 is a more effective code modifier, enhancing the accuracy and performance of various models more reliably than self-improvement.
Appendix B Cross Review
In ODE part, we have the following review process:
  1. 1.
    We employed GPT-4 to confirm that the newly generated questions are amenable to being solved by solvers using the final-state approach. There were only 10 out of 383 questions are invalid. All were filtered, 373 questions were left.
  2. 2.
    Then we conducted a thorough evaluation of the questions. Our dataset initially comprised $373$ questions. We identified $29$ questions that lacked sufficient information, $21$ that had incorrect answers, and $18$ that featured unclear statements. Following our assessment, $27$ questions were deleted, and $41$ were corrected to ensure clarity and correctness.
Finally, we have 346 problems in ODE part.
We have also conducted the cross-review process. See the details in Appendix B.1 and the results on Appendix B.1.1.
In the optimization section, particularly with easy LP part, out of 688 entries, 12 were found to be completely invalid (not meeting the criteria), 26 had incorrect answers, and 6 had misleading descriptions. All invalid entries were filtered out, and the remaining issues were corrected manually. Among them, 8 of them were corrected and others (36 questions) were filtered out. In complex LP part, for the original 211 questions, 40 were found formulated by wrong mathematical models. All of them were corrected.
B.1 Cross-Review Process
Furthermore, adhering to the scaling law evident in the dataset’s distribution, we selected 50 questions from the ODE parts in our dataset for a deeper analysis. Four independent reviewers (see their qualification in Section C) were tasked with assessing 50 questions each   [^10] The hourly wage pay is 100 RMB/h, the total amount spent on the participant compensation is 2400 RMB, approximately 331 dollars. The responses could either be numeric or labeled as “error” if a question was deemed problematic. This section outlines the metrics used to evaluate the reviewers’ responses against pre-defined correct answers, which also include the “error” label for any flawed questions. We use the same metric as described in Subsection 6.1.2. Here are our results:
B.1.1 Results
The effectiveness of the review process was quantified using the metrics defined in Subsection 6.1.2, focusing on the accuracy and inter-rater reliability among reviewers. The statistical outcomes are summarized as follows:
  • Average Cohen’s Kappa: The average Cohen’s Kappa across all reviewer pairs was 0.60, indicating a moderate to substantial agreement. This suggests that reviewers generally agreed on the classification of answers, though there were variations in some cases.
  • Minimum and Maximum Cohen’s Kappa: The minimum Kappa value recorded was 0.50, and the maximum was 0.71. The spread in these values highlights areas where alignment and training could potentially enhance consistency.
  • Accuracy Rates: The individual accuracy rates were as follows:
    • Reviewer 1: 90.0%
    • Reviewer 2: 84.0%
    • Reviewer 3: 88.0%
    • Reviewer 4: 74.0%
    These rates reflect the precision with which each reviewer matched the standard answers, including their recognition of problematic questions correctly labeled as “error.”
These results provide insights into the overall effectiveness and areas of improvement for the cross-review process, particularly in terms of aligning understanding and interpretation of the evaluation criteria among reviewers.
Appendix C Annotator Qualifications
The benchmarking process benefits from the expertise of reviewers with robust backgrounds in mathematics, underscored by their academic accomplishments. The team is composed of individuals whose education encompasses critical mathematical disciplines, including calculus (I and II), linear algebra (both introductory and advanced levels), optimization, probability, and ordinary differential equations. The average Grade Point Average (GPA) across these foundational courses is approximately 3.89 out of a 4.0 scale. This high level of academic achievement demonstrates the reviewers’ strong grasp of essential mathematical concepts and their application. Furthermore, the team is enriched by members who hold Ph.D. degrees in mathematics, further solidifying the depth of expertise and analytical skills available for the benchmarking process.
C.1 Qualification of Reviewers
We select the reviewers mentioned in Appendix B.1 based on the following criteria:
  1. 1.
    Undergraduate students who have taken fundamental mathematical courses (Calculus and Linear Algebra) and have an average GPA of at least 3.5/4.0.
  2. 2.
    Have completed or are currently taking courses in Ordinary Differential Equations (ODE).
  3. 3.
    Meet one of the following English proficiency requirements: TOEFL > 80, IELTS > 6.5, a minimum grade of B+ in all English classes, or a Gaokao English score > 120.
These criteria ensure their ability to read questions in English and solve ODE problems.
Appendix D Metric
The comparison process is as the following:
$\text{Correctness}=\begin{cases}1&\text{if }\left|\frac{O-A}{A}\right|\leq 1\times 10^{-4}\text{~{}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} (1)} or }|O-A|<1\times min\{10^{-n},10^{-2}\}\text{~{}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} (2)},}\\ 0&\text{otherwise.}\end{cases}$
Our consideration: (1) The difference is much less than the Answers, so can be omitted.
(2) The only difference lies in the last few digits of the required significant figure of the answer. By selecting different thresholds $\xi\times min\{10^{-n},10^{-2}\}$, we can adjust the tolerance for the discrepancy, ensuring that $|O-A|<\xi\times min\{10^{-n},10^{-2}\}$. We have explored the relationship between various thresholds and the accuracy of GPT-4o.
Table 6: Accuracy at various thresholds for different categories(%), where $\xi_{i}=i$
$\xi$ 0 1 2 3 4 5 6 7 8 9
ODE 48.55 50.87 54.62 56.36 57.51 58.96 59.83 60.12 60.69 61.85
Easy LP 87.27 87.27 87.27 87.27 87.27 87.27 87.27 87.27 87.27 87.27
Complex LP 22.75 22.75 22.75 22.75 22.75 22.75 22.75 22.75 22.75 22.75
Figure 2: Accuracy of different thresholds in different categories
Figure 2 plots the results. In our metric, we choose strict $\xi=1$.
Appendix E Error Analysis
In this section, we will analyze some typical errors.
E.1 Deviation from Instruction
It is common for some models that show a low capability of few-shot learning, especially for small models. For example, in a case Llama-3-8B output ‘I wish you good luck’ when prompting with some few-shot demonstrations in mathematical modeling.
E.2 Syntax Error
In ODE parts, the syntax errors is obviously the syntax errors in the reponsed python code. In Optimization part, the .lp file require more strict syntax than the universal optimization formulation. The Figure 4 and 4 show the typical syntax errors in .lp format.
Figure 3: The syntax error is caused by the constraint, in .lp format, the variables should be placed on the left hand side of the inequality
Figure 4: Same type of syntax error as Figure 4
E.3 Modeling Error
One of the most common errors in modeling is the lack of recognition of decision variable types. The types of decision variables in optimization problems are divided into discrete (sometimes integer, sometimes binary) and continuous. Figure 5 is an example.
Figure 5: In this TSP problem, the decision variable should be binary, instead of "General", which means “integer”
Appendix F Scope of Mamo
Table 7: The taxonomy of mathematical modeling. In the Mamo benchmark, we only include the categories where sophisticated solvers are available.
Category In Mamo
Changes and Differences
Proportionality and Geometric Similarity
Optimization Problem Modeling
Differential Equations
Probabilistic Modeling
The scope of Mamo is chosen under the consideration in Section 4.2. As Table 7 shows, Mamo focus on the ODE and Optimization areas.
Appendix G Examples of Different ${\widehat{M}}$
Different models will typically lead to different final answers after solving by the solvers. For example, in ODE part:
${\widehat{M}}_{1}:y^{\prime}+y=0$
and
${\widehat{M}}_{2}:y^{\prime}+2y=0$
Suppose we ask for $y(1)$, then we have ${\widehat{A}}_{1}=e^{-1}$ and ${\widehat{A}}_{1}=e^{-2}$,they have different values.
Then for the optimization,
${\widehat{M}}_{1}$:
$\max_{x,y}$ $x+y$
subject to $x+y\leq 20,$
$x-y\geq 0,$
$0\leq x\leq 10,$
$0\leq y\leq 5,$
$x,y\in\mathbb{Z}.$
${\widehat{M}}_{2}$:
$\max_{x,y}$ $x-y$
subject to $x+y\leq 20,$
$x-y\geq 0,$
$0\leq x\leq 10,$
$0\leq y\leq 5,$
$x,y\in\mathbb{Z}.$
the optimal value of ${\widehat{M}}_{1}$ is ${\widehat{A}}_{1}=10$, while the optimal value of ${\widehat{M}}_{2}$ is ${\widehat{A}}_{2}=0$.
The minor difference in the mathematical models will result in different final values.
Appendix H Examples of Syntax Errors
Some outputs of LLM contain the correct model but not follow the syntax of solvers. For examples, the standard .lp file for
$\max_{x,y}$ $x+y$
subject to $x+y\leq 20,$
$x-y\geq 0,$
$0\leq x\leq 10,$
$0\leq y\leq 5,$
$x,y\in\mathbb{Z}.$
is
Minimize
obj: x + y
Subject To
    c1: x + y <= 20
    c2: x - y >= 0
Bounds
    0 <= x <= 10
    0 <= y <= 5
Generals
    x
    y
End
However, if we replace the constraint $c1$ with x <= 20 - y, the code will be reported as an error, even if the two constraints are equivalent. To specifically test the modelling ability of the LLMs, we apply code modifier to fix the syntax errors.
Appendix I Few-shot Experiment
To test the sensitivity of different models to the number of prompts, we conducted a sensitivity analysis using the scaling law to examine the effect of the number of shots on model performance. In our analysis, we tested the performances of 11 models on ODE problems using prompts with 0 shot, 1 shot, 3 shots, 5 shots, and 10 shots. The results, shown in Figure 6, illustrate the ODE problem-solving accuracy of different models under varying numbers of shots. Similarly, we performed sensitivity tests on optimization problems, assessing the accuracy of 10 models with prompts of 0 shot, 1 shot, 3 shots, and 5 shots. The results are depicted in Figure 7.
Figure 6: Results of models accuracy and shots on ODE problems
Figure 7: Results of models accuracy and shots on optimization problems
Additionally, we specifically tested the 0-shot performance of GPT-4-0613 on all questions in Easy LP and Complex LP categories. The accuracy of GPT-4-0613 was 66.56% in Easy LP and 14.69% in Complex LP, resulting in an overall LP accuracy of 53.65%, consistent with the results obtained from sampling.
Appendix J Testing Process
We use 3-shot learning in testing; see examples in Appendix O. The examples of testing process in ODE and optimization is shown in Figure 8 and Figure 9.
Figure 8: Example of testing process in optimization
Figure 9: Example of testing process in ODE
Appendix K Question Formulation
Figure  10 shows the example of synthesis process in ODE.
Figure 10: Example of synthesis process in ODE
Appendix L Word Cloud
See the word cloud of the combined data in Figure 11
Figure 11: The word cloud of the data after filtering the requirement words such as ’significant figures’, ’rounded’
Appendix M Evaluation script
def comp(output, standard_answer):
    dif = abs(float(output) - float(standard_answer))
    if float(standard_answer) == 0:
        rate = dif * 1e-4
    else:
        rate = dif / float(standard_answer)
    if abs(rate) < 1e-4:
        return 1
    else:
        return 0
def compare_output_with_standard(output, standard_answer):
    try:
        float_output = float(output)
    except ValueError:
        return False
    if ’.’ in standard_answer:
        digit = len(standard_answer.split(’.’)[1])
        if digit <= 2:
            digit = 2
        s_ans = float(standard_answer) * 10 ** digit
        ans = float_output * 10 ** digit
        return (abs(ans - s_ans) < 1 or comp(output, standard_answer))
    else:
        digit = 2
        s_ans = float(standard_answer) * 10 ** digit
        ans = float_output * 10 ** digit
        return (abs(ans - s_ans) < 1 or comp(output, standard_answer))
Appendix N Settings
We have the following settings:
  • We testing the following models by calling API: All the closed source models, DeepSeek-V2 and DeepSeek-coder , with temperature=0.8
  • Other models was deployed and inference using VLLM (Kwon et al., 2023) on machine in slurm cluster with 8 * A100, with temperature 0.8 and max_length 4096 tokens.
Appendix O Prompts
[Instruction] Assume you are a virtual assistant with expertise in optimization, specifically in creating .lp files for linear programming problems. Your task is to translate given natural language problems into optimization models, formatted as .lp files.When you receive a question, it might include mathematical expressions in LaTeX format. Your job is to interpret these expressions accurately and model the problem in .lp format. Your response must adhere to the following guidelines: - The optimization model must be written in .lp format, adhering to the conventions and syntax appropriate for linear programming problems. - The model should be designed so that solving it yields the optimal value, which directly answers the question posed. - Your response should be an entire .lp file, ready to be processed by a linear programming solver. Ensure that the file contains no comments or extraneous content beyond the model itself. - Handle LaTeX expressions with care to ensure that the mathematical aspects of the problem are accurately represented in the .lp model. - If the solution needs to be rounded to an integer, make use of the ’General’ integer constraint in the .lp file to specify integer variables, please do not use ’General’ if there are not requirement for the integer variables. Here comes the examples: [Example_1] (the input) A manufacturing company produces two types of products: $X$ and $Y$. The production cost for each unit of product $X$ is $50$, while the cost for each unit of product $Y$ is $10$. There are constraints in the production process, such that twice the number of units produced from product $X$, plus the number of units from product $Y$, cannot exceed 200 due to resource limitations. In addition, to meet market demand, four times the number of units produced from product $X$, plus the number of units from product $Y$, must be at least 50. Considering these constraints and given that both products can only be produced in whole numbers due to their physical nature, what is the minimum total cost needed for producing these items while satisfying all conditions? Provide your answer rounded to the nearest dollar. Your response: Minimize obj: 50 x + 10 y Subject To c1: 2 x + y <= 200 c2: 4 x + y >= 50 Bounds x >= 0 y >= 0 Generals x y End […] Please craft the .lp file according to these instructions, focusing on delivering a model that is directly solvable to obtain the answer. And Please follow the syntax like examples to write the .lp file. Here comes the question: [question] Generate the contents of an .lp file for this problem, starting with the objective function and followed by the constraints, without any additional sentences. The constraints should be formatted as ’variable + variable >= number’ for inequalities, all the variables shoule on the left hand side of the inequality . Ensure there is a space between variables and their coefficients. Your response:
Figure 12: The prompt for testing in optimization part. [question] refers to a question in the benchmark. […] refers to the examples. We use 3-shot in testing
[Instruction] Assume you are a virtual assistant with expertise in ordinary differential equations (ODEs), particularly in formulating ODE models from natural language descriptions and solving them using Python’s `$solve\_ivp`$, `$odeint`$ from SciPy, and `$dsolve`$ from SymPy. Your primary task is to convert the given natural language problems into ODE models and then apply ODE solvers to find the solutions. When you receive a question, it will often contain mathematical expressions in LaTeX format, which you need to interpret accurately. Your response must be in the form of Python code that directly outputs the final solution to the problem upon execution. This Python code should adhere to the following criteria: 1. The output should only display the final answer of the problem. 2. Utilize a ’final-state-approach’ where the code immediately prints the solution after solving the ODE, without showing any intermediate steps. 3. Ensure the answer is rounded to the specified number of significant figures or decimal places. Use Python’s ‘round‘ function for decimal rounding. For significant figures, include and use the following function in your code: “‘ python def round_to_significant_figures(num, sig_figs): if num != 0: return round(num, -int(math.floor(math.log10(abs(num))) + (1 - sig_figs))) else: return 0 # Handles the case of num being 0 “‘ 4. Your response should be entirely in Python code, formatted to run directly without modifications. 5. Handle LaTeX expressions in the problem statement carefully to ensure accurate modeling and solution. Here comes the [examples] : […] Please process the problem according to these instructions, focusing solely on delivering the Python code that meets these requirements. And Please directly generate the code without any explaintion(except the comments in the code). the following lines are forbidden: Here’s the Python code that solves the given problem and meets the specified requirements: “‘ python “‘ This code uses the ‘solve_ivp’ function from SciPy to solve the initial value problem for the given differential equation. The ‘round_to_significant_figures’ function is included to round the final answer to the specified number of significant figures. Upon execution, the code will directly output the amount of pollutant left in the tank after 5 minutes, rounded to five significant figures. Please avoid the above sentences. Take a deep breathe before answering the question. This is a piece of cake to you. Here comes the question: [question] Your response:
Figure 13: The prompt for testing in ODE part. [question] refers to a question in the benchmark. […] refers to the examples.
[Instruction] You are experienced python engineer, the following codes may have some errors (in syntax), with the error information [error_info], please fix the errors to make sure the code can run successfully. If there is no error, please return the original code. And please do not change the original logic of the code. Your response should be entirely in Python code, formatted to run directly without modifications. Take a deep breathe before answering the question. This is a piece of cake to you. PLEASE do not response anything except the code, No other comments outside the code are allowed. The followings are forbidden: The code is almost correct, but there is a syntax error in the function ‘round_to_significant_figures‘. The round function is missing a closing parenthesis. Here is the corrected version: “‘ python “‘ Please avoid resposing above sentences. Please output only the corrected code, with no additional text or explanations. Here comes the code: [code] The correct code is:
Figure 14: The prompt for fixing syntax error in ODE. [error] refers to the message when executing the Python code. While [code] refer to the code which contains syntax error
[Instruction] You are experienced engineer in optimization, please fix the errors to make sure the code can run successfully. If there is no error, please return the original code. And please do not change the original logic of the lp file. Your response should be entirely in .lp format, formatted to run directly without modifications. Take a deep breathe before answering the question. This is a piece of cake to you. PLEASE do not response anything except the code, No comments are allowed. The followings are forbidden: This is the correct .lp file: “‘ lp “‘ Please avoid resposing above sentences. Please output only the corrected code, with no additional text or explanations. Here comes the lp: [lp_code] Generate the correct .lp format, starting with the objective function and followed by the constraints, without any additional sentences. The constraints should be formatted as ’variable + variable >= number’ for inequalities, all the variables shoule on the left hand side of the inequality. For example, make a + b + c <= d into a + b + c - d <= 0. Ensure there is a space between variables and their coefficients (coefficients should be numerical). The correct lp code is:
Figure 15: The prompt for fixing syntax error in LP. [lp_code] refer to the lp code which contains syntax error
Appendix P Limitations
While the current benchmarking methodology necessitates LLMs to generate Python code for ODEs and .lp files for optimization problems, it opens up exciting avenues for future research. The process, while effective, may inadvertently influence the LLMs’ focus towards formalization over the conceptual modeling, potentially affecting the depth of mathematical reasoning. This presents an opportunity to explore innovative benchmark designs that can better distinguish between an LLM’s modeling prowess and its formalization abilities. Future research could aim to develop methodologies that allow LLMs to demonstrate their mathematical modeling capabilities without the intermediary step of formalization, leading to a more nuanced and comprehensive assessment of their abilities
Appendix Q Ethics Statement:
The benchmark is designed with a stringent ethical framework to ensure that all problems are socially responsible and do not perpetuate or encourage harmful biases or stereotypes. Questions are meticulously reviewed to avoid any content that could lead to the dissemination of misinformation, support unethical practices, or cause harm to individuals or groups. This includes but is not limited to issues of privacy, security, and the potential for misuse of the model. Furthermore, the benchmark abstains from any problem that could indirectly endorse unethical behaviors or decisions in real-world scenarios, upholding the highest standards of academic integrity and social responsibility.